Suppose the value of a stock varies each day from $12 to $27 with a uniform distribution. Find ?
Answer #1

Let X be a Uniform random variable.

Here, X ~ Uniform ( a = $12, b = $ 27)

Mean and standard deviation of Uniform distribution as follows,

\text { Mean } = \mu = \frac {a+b}{2} =\frac{ 12+27}{2}=\frac{ 39}{2}= 19.5

\mu = 19.5

\text{ Standard deviation } (\sigma) = \sqrt { \frac { (b-a)^2}{12}}= \sqrt{ \frac{ (27-12)^2}{12}}= 4.33012702

\sigma= 4.3301